Post Card Depot, an large retailer of post cards, orders 5,495,510 post cards per year from its manufacturer. Post Card Depot plans on ordering post card 20 times over the next year. Post Card Depot receives the same number of post cards each time it orders. The carrying cost is $0.30 per post card per year. The ordering cost is $407 per order.

What is the annual carrying costs of post card inventory
Answer #1

Annual carrying costs = (Q/2) × C

where, Q = the inventory order size in units

C = carrying cost per unit

Q (the inventory order size in units) = 5,495,510 / 20 = 274,775.5 post cards

Annual carrying costs = (274,775.5/2) × $0.30

Annual carrying costs = $41,216.33